Please explain and show work for each

6. Consider the decay scheme of xenon-133. a. Comp
Answer #1

6 a) maximum energy of the three beta particles = 346.4+266.8+43.6

= 656.8N KeV

B) Average energy = 656.8/3 = 218.9 KeV

C) Energies of six gamma particles = 79.6+81+160.6+223.2+302.8+383.8

= 1231 KeV = 1.231 MeV